# A Lightweight 3D Anomaly Detection Method with Rotationally Invariant Features

### v1

@@ Abstract
3D anomaly detection (AD) is a crucial task in computer vision, aiming to identify anomalous points or regions from point cloud data. However, existing methods may encounter challenges when handling point clouds with changes in orientation and position because the resulting features may vary significantly. To address this problem, we propose a novel Rotationally Invariant Features (RIF) framework for 3D AD. Firstly, to remove the adverse effect of variations on point cloud data, we develop a Point Coordinate Mapping (PCM) technique, which maps each point into a rotationally invariant space to maintain consistency of representation. Then, to learn robust and discriminative features, we design a lightweight Convolutional Transform Feature Network (CTF-Net) to extract rotationally invariant features for the memory bank. To improve the ability of the feature extractor, we introduce the idea of transfer learning to pre-train the feature extractor with 3D data augmentation. Experimental results show that the proposed method achieves the advanced performance on the Anomaly-ShapeNet dataset, with an average P-AUROC improvement of 17.7%, and also gains the best performance on the Real3D-AD dataset, with an average P-AUROC improvement of 1.6%. The strong generalization ability of RIF has been verified by combining it with traditional feature extraction methods on anomaly detection tasks, demonstrating great potential for industrial applications. Source code is available at https://github.com/hzzzzzhappy/RIF.
keywords:
3D anomaly detection, point cloud, rotationally invariant, convolution, industrial anomaly detection.
journal: Pattern Recognition
[…]
1 Introduction
Three Dimension anomaly detection (3D-AD) [zhu2024towards, li2023scalable3danomalydetection, AdaCLIP] has attracted increasing attention from the computer vision community due to its widespread application in inspecting high-precision industrial products. It is dedicated to identifying anomalous points or regions that deviate from given 3D point cloud data. In real-world industrial applications, we often face a scenario where it is easy to collect normal samples but time-consuming and costly to acquire defective samples, such that unsupervised methods that only learn from normal samples are widely used.
@@ 1 Introduction
Feature embedding-based methods use pre-trained models to extract features of normal samples for a memory bank, with anomalies detected by comparing test sample features to those in the bank. Some methods such as M3DM [Wang2023multimodal] have achieved on the datasets like MvTec3D-AD [Bergmann_2022] using memory bank methods. However, they face challenges on unregistered datasets such as Real3D-AD [Liu2023real3d] and Anomaly-ShapeNet [li2023scalable3danomalydetection]. Reg3D-AD [Liu2023real3d] used PointMAE [Pang2022pointmae] for feature representation based on the registered point clouds, and Group3AD [zhu2024towards] further used contrast learning to improve feature differentiation ability. On the other hand, reconstruction-based methods encode point cloud data into a latent space and decode it in its original form, with high reconstruction errors indicating anomalies. IMRNet [li2023scalable3danomalydetection] improved anomaly detection by using geometric-preserving downsampling and random masking. R3D-AD [zhou2025r3dad] utilized PointNet [Qi_2017_CVPR] and a diffusion process to progressively restore the point cloud. These methods have shown promising results on the Anomaly-ShapeNet [li2023scalable3danomalydetection] and Real3D-AD [Liu2023real3d] datasets.
Industrial 3D AD is particularly difficult because products on the production line may appear in different poses [cheng2025highresolution3danomalydetection]. Consequently, the same product may have point cloud data with variations in orientation and position. Although a rotating point cloud does not alter its geometric structure, it changes the coordinate representation corresponding to the structure, resulting in a significant difference in the output of the feature extraction network. Reg3D-AD [Liu2023real3d] aligns test samples with a standard template via registration to reduce data variation effects, but this method is computationally intensive and risks registration errors, particularly with significant train-test gaps. SplatPose [Kruse2024spkatpose] reconstructs 3D objects via Gaussian splatting to enhance pose robustness; however, its point cloud precision relies on image quality and view count, and it requires test image-to-3D alignment. This inspires us to find a unique representation of point cloud data to maintain rotation invariance. Therefore, our method proposed to map point clouds into a rotational invariance space through coordinate transformation, enabling the generation of a unique feature representation for point clouds with different rotations and translations. The differences between our solution and previous methods are shown in Figure 1.
@@
The rest of this paper is organized as follows. Section 2 recalls some related work. Section 3 details our proposed method. In Section 4, extensive experiments are conducted to validate the proposed method. Finally, Section 5 draws the conclusion of the paper.
2 Related Work
2.1 2D Anomaly Detection
@@ 2 Related Work
2D anomaly detection [Zhao2023omnial, LEE2026111820, LIU2026112198, LI2024110258, MIAO2026112163, tang2025mpr] is an important vision task that aims to detect and localize anomalies from images. Existing 2D AD methods are mainly classified into two categories: generative and discriminative methods [PangSC2021, 10034849]. Generative models such as Autocoder [Gong2019memae], GAN [Liang2023ocrgan], and Diffusion [he2024diad] are employed to learn the feature representation or distribution of normal samples, and anomalies are detected by performing image reconstruction [ZAVRTANIK2021reconstruction] or matching with a memory bank  [roth2022totalrecallindustrialanomaly]. Recent methods, such as DRAEM [Zavrtanik2021DRAEM], PatchCore [roth2022totalrecallindustrialanomaly] and CRAD [Lee2024crad] have yielded impressive results. Discriminative methods aim to train a supervised model to discriminate anomalous and normal images, which usually requires a certain number of anomalous samples annotated by humans or generated by data augmentation. Some methods like DRA [Ding2022catch], BGAD [Zhang2023prototypical], and AHL [Zhu2024openset] have shown the benefits of supervised information to improve 2D anomaly detection performance. Although promising results, applying 2D anomaly detection methods to 3D point clouds directly may face significant challenges.
2.2 3D Anomaly Detection
@@
3D anomaly detection intends to identify and locate anomalous points or regions within 3D point cloud data. Existing methods for 3D AD can be grouped into two main categories: feature embedding-based methods and reconstruction-based methods. Feature embedding-based methods use pre-trained models to extract normal features for a memory bank, detecting anomalies by comparing test features against stored entries [10702559, ZHANG2025111360, 1asda]. Reg3D-AD [Liu2023real3d, 10260338] uses PointMAE [Pang2022pointmae] to extract global geometric and local coordinate features post-registration, matching both during inference for anomaly scoring. Group3AD [zhu2024towards] clusters groups into uniformly compact structures, constructing group-level features for the memory bank. Reconstruction-based methods aim to project point cloud data into a high-dimensional latent space through encoding processes, subsequently reconstructing the original input via decoding, with points demonstrating elevated reconstruction errors being identified as anomalies. IMRNet [li2023scalable3danomalydetection] improved PointMAE by geometric-preserving downsampling and random masking to enhance reconstruction ability for anomaly detection. R3DAD [zhou2025r3dad] employed PointNet to progressively restore full-mask point clouds by performing the diffusion, aiding the model in accurately identifying abnormal regions. Additionally, some methods integrate 2D images with 3D point clouds to perform RGB-D AD. M3DM [Wang2023multimodal] aligns PointTransformer (3D) and ViT (2D) features via contrastive learning. CPMF [cao2023complementarypseudomultimodalfeature] combines local 3D geometry with 2D semantics from multi-view images. Looking3D [Bhunia2024look] integrates texture-enhanced 3D data with cross-modality matching.
@@
3 Approach
The core of our RIF framework is to learn rotationally invariant features to address the diversity of samples in orientation, position, and posture. Here, we propose a simple yet efficient solution for 3D AD, which is shown in Figure 2. We will elaborate on its details in the following sections. To improve the understanding, the main symbols in the paper and their explanations are reported in Table 1.
@@ 3 Approach
Table 1: Main symbols and their explanations in the paper.
3.1 Point Coordinate Mapping for Rotational Invariance
@@
  • 3.
    $\mathbf{u}_{3}=\mathbf{p}_{\text{near}}-\mathbf{c}$, where $\mathbf{p}_{\text{near}}$ is the nearest point to the centroid that makes the rank of the three selected vectors equal to 3, i.e., $Rank([\mathbf{u}_{1};$ $\mathbf{u}_{2};\mathbf{u}_{3}])=3$.
@@
3.2 Pre-trained Feature Extractor with Spatial 3D Data Augmentation
$\textbf{P}_{(n,3)}^{\prime}=Z(J(S(\textbf{P}_{(n,3)}))),$ (2)
3.3 Convolutional Transform Feature Network for Anomaly Detection
$\textbf{p}_{i}$ $=FPS(PCM(\textbf{P}_{(n,3)})),$ (6)
$\textbf{G}_{i}$ $=KNN(\textbf{p}_{i}),$ (7)
$\textbf{Mem}_{i}=CTF\text{-}Net(\textbf{G}_{i}).$ (8)
$\textbf{S}_{i}=\parallel\textbf{Mem}_{near}-\textbf{F}_{i}\parallel,$ (9)
4 Experiments
@@ 4.1 Experimental Setup
4.2 Experimental Results and Analysis
To show the validity of our method, we conducted comparison experiments on Anomaly-ShapeNet and Real3D-AD, and the performance in terms of P-AUROC is shown in Tables 2 and  3, respectively. Prior work has demonstrated that rotation-invariant representation methods often exhibit diminished performance on general downstream tasks [ao2021spinnetlearninggeneralsurface, HE2025107223]. In contrast, our proposed approach successfully reconciles this trade-off by achieving state-of-the-art AD performance while preserving robust rotation-invariant properties.
@@ 4.2 Experimental Results and Analysis
Table 2: P-AUROC performance of different methods on Anomaly-ShapeNet across 40 categories, where the best and second-place results are boldfaced and underlined, respectively.
@@
Table 3: P-AUROC performance of different methods on Real3D-AD across 12 categories, where the best and second-place results are boldfaced and underlined, respectively.
1) Comparison on Anomaly-ShapeNet. As shown in Table 2, the proposed method achieved the best average P-AUROC of 84.5%, which is improved over the second-place method Reg3D-AD by 17.7% and R3D-AD by 18.8% (author report: P-AUROC is 65.7%). Notably, for the “cap0” and “jar” categories, the P-AUROC values attained by our method are 96.8% and 96.7%, respectively, whereas Reg3D-AD only achieved 73.0% and 76.5%. This significant improvement could potentially be attributed to the ability of the proposed method to adapt samples with variations using a unified feature representation, thereby enhancing the discriminative power of the extracted features, particularly in sparse point clouds. Furthermore, the utilization of a large number of points for each group ($K=512$) may have contributed to the superior anomaly detection performance in such sparse point cloud scenario.
2) Comparison on Real3D-AD. Table 3 shows the pixel-level detection performance of all methods on Real3D-AD. Our method achieved 78.3% P-AUROC on average. Among all selected methods, ours performs the best in terms of P-AUROC, surpassing the second-place method by 1.6%. Particularly, for the “Candybar” and “Diamond” categories, the P-AUROC values of our method are 91.3% and 97.5%, respectively. The reason for this may be attributed to different point clouds sharing similar structures may have large variations in orientation and position, leading to large differences in the extracted features, while our method exhibits great robustness in capturing such variations compared to other methods. Moreover, for large-scale point clouds like Real3D-AD, our method obtained superior representation and performance with fewer groups ($G=512$) compared to Reg3D-AD ($G=16384$) and IMRNet (downsampled to 8192 points). This suggests that our CTF-Net could extract better features for 3D anomaly detection.
4.3 Ablation Studies and Parameter Analysis
To evaluate the importance of key components and the effect of parameters on our method, we conducted ablation studies and parameter sensitivity analysis on Anomaly-ShapeNet, and the results are shown in Table 4.
Method P-AUROC O-AUROC P-AUPRO O-AUPRO FPS Memory
@@ 4.3 Ablation Studies and Parameter Analysis
OursCCB=1 0.7263 0.6210 0.1364 0.6824 0.42 19.11%
OursCCB=2 0.7588 0.6223 0.1410 0.6963 0.39 19.94%
OursCCB=3 0.8001 0.6327 0.1684 0.7163 0.36 21.92%
OursCCB=4 0.8253 0.6609 0.2029 0.7432 0.35 23.33%
OursCCB=5 0.8446 0.6893 0.2124 0.7480 0.31 25.85%
OursG=64,K=512 0.6694 0.5960 0.0553 0.6655 0.50 16.34%
OursG=128,K=512 0.7456 0.6070 0.1021 0.6826 0.43 19.68%
OursG=256,K=512 0.8117 0.6692 0.1639 0.7318 0.36 20.43%
OursG=512,K=64 0.7919 0.6127 0.1193 0.6788 1.54 12.23%
OursG=512,K=128 0.7782 0.6297 0.0973 0.6809 1.02 14.42%
OursG=512,K=256 0.8215 0.6487 0.1532 0.7087 0.68 16.84%
OursG=512,K=512 0.8446 0.6893 0.2124 0.7480 0.31 25.85%
Table 4: Results of ablation and parameter sensitivity experiments, where w/o PCM, w/o S3DA, and w/o Trans. mean the proposed method without PCM, S3DA, and transformation matrix, respectively, replace RANSAC, replace PCA, and replace CCB mean replacing our PCM with RANSAC or PCA and replacing CCB with 1D convolution, respectively; Parameter CCB denotes the number of blocks in the feature extractor CTF-Net, $G$ and $K$ represent the number of groups and the number of points within each group, respectively. Frames Per Second (FPS) and Memory (%) stand for the average inference speed and memory consumption of our CTF-Net with an RTX3090, respectively.
@@
Our CTF-Net has only 4.1M parameters, much less than other methods such as PointMLP (12.6M), and memory consumption is only 26% on an RTX3090. In addition, our method achieved an FPS of 0.33, outperforming other methods such as Reg3D-AD (FPS of 0.13).
The complexity of environments and equipment instability may result in scanned point clouds containing noise in real-world scenarios. To evaluate the performance of our model with noise, we conducted comparative experiments on point clouds with different standard Gaussian noises under one-shot conditions. For the fairness comparison, all pre-training was performed on ModelNet40 [sun2022modelnet40]. The results for Gaussian noise with a standard deviation of 0.002 and 0.003 are shown in Tables 5 and 6, respectively. It is observed that the performance decreases only slightly as the standard deviation of the noise increases, and is more noise-resistant compared to other approaches, showing better anomaly detection performance(P-AUROC at standard deviations of 0.02 and 0.03 was 9.8% and 4.8% higher than the baseline model Reg3D-AD [Liu2023real3d], respectively). This may be attributed to the ability of rotationally invariant features to capture features more robustly under noisy conditions. These empirical results demonstrate the robustness of our method in detecting anomalies from noisy point clouds.
@@
Table 5: One-Shot performance of different methods on Real3D-AD across 12 categories under Gaussian noise with a standard deviation of 0.002, where the best results are boldfaced.
@@
Table 6: One-Shot performance of different methods on Real3D-AD across 12 categories under Gaussian noise with a standard deviation of 0.003, where the best results are boldfaced.
4.4 Robustness Experimental Results to Noisy Data.
4.5 Generalization of Our Framework RIF
To examine the RIF’s generalization, we conducted extensibility experiments on Anomaly-ShapeNet. The feature extractor CTF-Net was replaced by other representative point cloud feature extractors, including PointNet [pointnet], PointMLP [pointmae], DGCNN [dgcnn], PointMAE [pointmae], Stratified-Transformer [stratifiedtransformer], and PointTransformer-v3 [pointtransformerv3], and some results are reported in Table 7.

### v2

@@ Abstract
3D anomaly detection (AD) is a crucial task in computer vision, aiming to identify anomalous points or regions from point cloud data. However, existing methods may encounter challenges when handling point clouds with changes in orientation and position because the resulting features may vary significantly. To address this problem, we propose a novel Rotationally Invariant Features (RIF) framework for 3D AD. Firstly, to remove the adverse effect of variations on point cloud data, we develop a Point Coordinate Mapping (PCM) technique, which maps each point into a rotationally invariant space to maintain consistency of representation. Then, to learn robust and discriminative features, we design a lightweight Convolutional Transform Feature Network (CTF-Net) to extract rotationally invariant features for the memory bank. To improve the ability of the feature extractor, we introduce the idea of transfer learning to pre-train the feature extractor with 3D data augmentation. Experimental results show that the proposed method achieves the advanced performance on the Anomaly-ShapeNet dataset, with an average P-AUROC improvement of 17.7%, and also gains the best performance on the Real3D-AD dataset, with an average P-AUROC improvement of 1.6%. The strong generalization ability of RIF has been verified by combining it with traditional feature extraction methods on anomaly detection tasks, demonstrating great potential for industrial applications.
I Introduction
Three Dimension anomaly detection (3D-AD) [zhu2024towards, li2023scalable3danomalydetection] has attracted increasing attention from the computer vision community due to its widespread application in inspecting high-precision industrial products. It is dedicated to identifying anomalous points or regions that deviate from given 3D point cloud data. In real-world industrial applications, we often face a scenario where it is easy to collect normal samples but time-consuming and costly to acquire defective samples, such that unsupervised methods that only learn from normal samples are widely used.
@@ I Introduction
Feature embedding-based methods use pre-trained models to extract features of normal samples for a memory bank, with anomalies detected by comparing test sample features to those in the bank. Some methods such as M3DM [Wang2023multimodal] have achieved on the datasets using memory bank methods. However, they face challenges on unregistered datasets such as Real3D-AD [Liu2023real3d] and Anomaly-ShapeNet [li2023scalable3danomalydetection]. Reg3D-AD [Liu2023real3d] used PointMAE [Pang2022pointmae] for feature representation based on the registered point clouds, and Group3AD [zhu2024towards] further used contrast learning to improve feature differentiation ability. On the other hand, reconstruction-based methods encode point cloud data into a latent space and decode it in its original form, with high reconstruction errors indicating anomalies. IMRNet [li2023scalable3danomalydetection] improved anomaly detection by using geometric-preserving downsampling and random masking. R3D-AD [zhou2025r3dad] utilized PointNet [Qi_2017_CVPR] and a diffusion process to progressively restore the point cloud. These methods have shown promising results on the Anomaly-ShapeNet [li2023scalable3danomalydetection] and Real3D-AD [Liu2023real3d] datasets.
Industrial 3D AD is particularly difficult because products on the production line may appear in different poses [cheng2025highresolution3danomalydetection]. Consequently, the same product may have point cloud data with variations in orientation and position. Although a rotating point cloud does not alter its geometric structure, it changes the coordinate representation corresponding to the structure, resulting in a significant difference in the output of the feature extraction network. Reg3D-AD [Liu2023real3d] aligns test samples with a standard template via registration to reduce data variation effects, but this method is computationally intensive and risks registration errors, particularly with significant train-test gaps. SplatPose [Kruse2024spkatpose] leverages novel view synthesis via Gaussian splatting to reconstruct 3D objects and enhance pose robustness; however, its point cloud fidelity still depends heavily on image quality, synthesized-view consistency, and view count, and it requires test-time image–3D alignment. These limitations motivate us to pursue a purely geometric and view-free representation that preserves rotation invariance directly on point clouds. Therefore, our method proposed to map point clouds into a rotational invariance space through coordinate transformation, enabling the generation of a unique feature representation for point clouds with different rotations and translations. The differences between our solution and previous methods are shown in Figure 1.
@@
The rest of this paper is organized as follows. Section II recalls some related work. Section III details our proposed method. In Section IV, extensive experiments are conducted to validate the proposed method. Finally, Section V draws the conclusion of the paper.
II Related Work
II-A 2D Anomaly Detection
@@ II Related Work
2D anomaly detection [Zhao2023omnial, LEE2026111820, LIU2026112198, LI2024110258, MIAO2026112163, tang2025mpr] is an important vision task that aims to detect and localize anomalies from images. Existing 2D AD methods are mainly classified into two categories: generative and discriminative methods [PangSC2021, 10034849]. Generative models such as Autocoder [Gong2019memae], GAN, and Diffusion are employed to learn the feature representation or distribution of normal samples, and anomalies are detected by performing image reconstruction [ZAVRTANIK2021reconstruction, luo2025exploring] or matching with a memory bank [roth2022totalrecallindustrialanomaly, luo2025ura]. Recent methods, such as DRAEM [Zavrtanik2021DRAEM], PatchCore [roth2022totalrecallindustrialanomaly] and CRAD have yielded impressive results. Discriminative methods aim to train a supervised model to discriminate anomalous and normal images, which usually requires a certain number of anomalous samples annotated by humans or generated by data augmentation. Some methods BGAD [Zhang2023prototypical], and AHL [Zhu2024openset] have shown the benefits of supervised information to improve 2D anomaly detection performance. Although promising results, applying 2D anomaly detection methods to 3D point clouds directly may face significant challenges.
II-B 3D Anomaly Detection
@@
3D anomaly detection intends to identify and locate anomalous points or regions within 3D point cloud data [3Dsurvey, liang2025taming]. Existing methods for 3D AD can be grouped into two main categories: feature embedding-based methods and reconstruction-based methods. Feature embedding-based methods use pre-trained models to extract normal features for a memory bank, detecting anomalies by comparing test features against stored entries [10702559, ZHANG2025111360, 1asda, 10898004]. Reg3D-AD [Liu2023real3d, 10260338] uses PointMAE [Pang2022pointmae] to extract global geometric and local coordinate features post-registration, matching both during inference for anomaly scoring. Group3AD [zhu2024towards] clusters groups into uniformly compact structures, constructing group-level features for the memory bank. Reconstruction-based methods aim to project point cloud data into a high-dimensional latent space through encoding processes, subsequently reconstructing the original input via decoding, with points demonstrating elevated reconstruction errors being identified as anomalies. PASDF [zheng2025bridging] learns high-quality continuous representations of the signed distance for each point and directly identifies outlier anomalies based on the signed distances. IMRNet [li2023scalable3danomalydetection] improved PointMAE by geometric-preserving downsampling and random masking to enhance reconstruction ability for anomaly detection. R3DAD [zhou2025r3dad] employed PointNet to progressively restore full-mask point clouds by performing the diffusion, aiding the model in accurately identifying abnormal regions. Additionally, some methods integrate 2D images with 3D point clouds to perform RGB-D AD. M3DM [Wang2023multimodal] aligns PointTransformer (3D) and ViT (2D) features via contrastive learning. CPMF [cao2023complementarypseudomultimodalfeature] combines local 3D geometry with 2D semantics from multi-view images. Looking3D [Bhunia2024look] integrates texture-enhanced 3D data with cross-modality matching.
@@
Traditional descriptors such as FPFH and multi-FPFH rely on statistical histograms of local geometric relations, providing only implicit rotation robustness. In contrast, PCM constructs an explicit rotation-invariant coordinate frame via deterministic geometric anchoring and orthogonalization, making its mathematical formulation fundamentally different from these histogram-based operators.
III Approach
The core of our RIF framework is to learn rotationally invariant features to address the diversity of samples in orientation, position, and posture. Here, we propose a simple yet efficient solution for 3D AD, which is shown in Figure 2. We will elaborate on its details in the following sections. To improve the understanding, the main symbols in the paper and their explanations are reported in Table I.
@@ III Approach
TABLE I: Main symbols and their explanations in the paper.
III-A Point Coordinate Mapping for Rotational Invariance
@@
  • $\mathbf{u}_{3}=\mathbf{p}_{\text{near}}-\mathbf{c}$, where $\mathbf{p}_{\text{near}}$ is the nearest point to the centroid that makes the rank of the three selected vectors equal to 3, i.e., $Rank([\mathbf{u}_{1};$ $\mathbf{u}_{2};\mathbf{u}_{3}])=3$;
  • If multiple farthest or nearest points satisfy $\|\mathbf{p}-\mathbf{c}\|=d_{\max}$ or $d_{\min}$, we replace them with their centroid to resolve distance ties, yielding a deterministic representative point and a stable PCM basis.
@@
III-B Pre-trained Feature Extractor with Spatial 3D Data Augmentation
$\textbf{P}_{(n,3)}^{\prime}=Z(J(S(\textbf{P}_{(n,3)}))),$ (2)
III-C Convolutional Transform Feature Network for Anomaly Detection
$\textbf{p}_{i}$ $=FPS(PCM(\textbf{P}_{(n,3)})),$ (6)
$\textbf{G}_{i}$ $=KNN(\textbf{p}_{i}),$ (7)
$\textbf{Mem}_{i}=CTF\text{-}Net(\textbf{G}_{i}).$ (8)
$\textbf{S}_{i}=\parallel\textbf{Mem}_{near}-\textbf{F}_{i}\parallel,$ (9)
@@ III-C Convolutional Transform Feature Network for Anomaly Detection
IV Experiments
@@ IV-A Experimental Setup
IV-B Experimental Results and Analysis
To show the validity of our method, we conducted comparison experiments on Anomaly-ShapeNet and Real3D-AD, and the performance in terms of P-AUROC is shown in Table II and Table III, respectively. Prior work has demonstrated that rotation-invariant representation methods often exhibit diminished performance on general downstream tasks. In contrast, our proposed approach successfully reconciles this trade-off by achieving state-of-the-art AD performance while preserving robust rotation-invariant properties.
@@ IV-B Experimental Results and Analysis
TABLE II: P-AUROC performance of different methods on Anomaly-ShapeNet across 40 categories.
@@
TABLE III: P-AUROC performance of different methods on Real3D-AD across 12 categories, where the best and second-place results are boldfaced and underlined, respectively.
1) Comparison on Anomaly-ShapeNet. As shown in Table 2, the proposed method achieved the best average P-AUROC of 84.5%, which is improved over the competitive approach Reg3D-AD by 17.7%, DUS-Net by 13.3%, and R3D-AD by 18.8%. Notably, for the “cap0” and “jar” categories, the P-AUROC values attained by our method are 96.8% and 96.7%, respectively, whereas Reg3D-AD only achieved 73.0% and 76.5%. This significant improvement could potentially be attributed to the ability of the proposed method to adapt samples with variations using a unified feature representation, thereby enhancing the discriminative power of the extracted features, particularly in sparse point clouds. Furthermore, the utilization of a large number of points for each group ($K=512$) may have contributed to the superior anomaly detection performance in such sparse point cloud scenario.
2) Comparison on Real3D-AD. Table III shows the pixel-level detection performance of all methods on Real3D-AD. Our method achieved 78.3% P-AUROC on average. Among all selected methods, ours performs the best in terms of P-AUROC, surpassing the second-place method by 1.6%. Particularly, for the “Candybar” and “Diamond” categories, the P-AUROC values of our method are 91.3% and 97.5%, respectively. The reason for this may be attributed to different point clouds sharing similar structures may have large variations in orientation and position, leading to large differences in the extracted features, while our method exhibits great robustness in capturing such variations compared to other methods. Moreover, for large-scale point clouds like Real3D-AD, our method obtained superior representation and performance with fewer groups ($G=512$) compared to Reg3D-AD ($G=16384$) and IMRNet (downsampled to 8192 points). This suggests that our CTF-Net could extract better features for 3D anomaly detection.
IV-C Ablation Studies and Parameter Analysis
To evaluate the importance of key components and the effect of parameters on our method, we conducted ablation studies and parameter sensitivity analysis on Anomaly-ShapeNet, and the results are shown in Table IV, where the performance, efficiency, and memory consumption, including both the percentage and the absolute values of memory usage, are reported.
Method P-AUROC O-AUROC P-AUPRO O-AUPRO FPS Memory(%) Memory(GB)
@@ IV-C Ablation Studies and Parameter Analysis
OursCCB=1 0.7263 0.6210 0.1364 0.6824 0.42 19.11% 4.59
OursCCB=2 0.7588 0.6223 0.1410 0.6963 0.39 19.94% 4.79
OursCCB=3 0.8001 0.6327 0.1684 0.7163 0.36 21.92% 5.26
OursCCB=4 0.8253 0.6609 0.2029 0.7432 0.35 23.33% 5.60
OursCCB=5 0.8446 0.6893 0.2124 0.7480 0.31 25.85% 6.20
OursG=64,K=512 0.6694 0.5960 0.0553 0.6655 0.50 16.34% 3.92
OursG=128,K=512 0.7456 0.6070 0.1021 0.6826 0.43 19.68% 4.72
OursG=256,K=512 0.8117 0.6692 0.1639 0.7318 0.36 20.43% 4.90
OursG=512,K=64 0.7919 0.6127 0.1193 0.6788 1.54 12.23% 2.94
OursG=512,K=128 0.7782 0.6297 0.0973 0.6809 1.02 14.42% 3.46
OursG=512,K=256 0.8215 0.6487 0.1532 0.7087 0.68 16.84% 4.04
OursG=512,K=512 0.8446 0.6893 0.2124 0.7480 0.31 25.85% 6.20
TABLE IV: Results of ablation and parameter sensitivity experiments, where w/o PCM, w/o S3DA, and w/o Trans. mean the proposed method without PCM, S3DA, and transformation matrix, respectively, replace RANSAC, replace PCA, and replace CCB mean replacing our PCM with RANSAC or PCA and replacing CCB with 1D convolution, respectively; Parameter CCB denotes the number of blocks in the feature extractor CTF-Net, $G$ and $K$ represent the number of groups and the number of points within each group, respectively. Frames Per Second (FPS) and Memory (% and GB) stand for the average inference speed and memory consumption of our CTF-Net with an RTX3090, respectively.
@@
Our CTF-Net contains only 4.1M parameters, significantly fewer than existing alternatives such as PointMLP (12.6M), and requires only 6.2 GB of GPU memory on an RTX3090, which is approximately 0.9 GB less than memory-bank–based methods such as Reg3D-AD (7.1 GB). In addition, our approach achieved an inference speed of 0.33 FPS, outperforming methods like Reg3D-AD, which reports only 0.13 FPS.
IV-D Robustness Experimental Results to Noisy Data.
The complexity of environments and equipment instability may result in scanned point clouds containing noise in real-world scenarios. To evaluate the performance of our model with noise, we conducted comparative experiments on point clouds with different standard Gaussian noises under one-shot conditions. For the fairness comparison, all pre-training was performed on ModelNet40 [sun2022modelnet40]. The results for Gaussian noise with a standard deviation of 0.002 and 0.003 are shown in Tables V and VI, respectively. It is observed that the performance decreases only slightly as the standard deviation of the noise increases, and is more noise-resistant compared to other approaches, showing better anomaly detection performance(P-AUROC at standard deviations of 0.02 and 0.03 was 9.8% and 4.8% higher than the baseline model Reg3D-AD [Liu2023real3d], respectively). This may be attributed to the ability of rotationally invariant features to capture features more robustly under noisy conditions. These empirical results demonstrate the robustness of our method in detecting anomalies from noisy point clouds.
@@
TABLE V: One-Shot performance of different methods on Real3D-AD across 12 categories under Gaussian noise with a standard deviation of 0.002, where the best results are boldfaced.
@@
TABLE VI: One-Shot performance of different methods on Real3D-AD across 12 categories under Gaussian noise with a standard deviation of 0.003, where the best results are boldfaced.
IV-E Generalization of Our Framework RIF
To examine the RIF’s generalization, we conducted extensibility experiments on Anomaly-ShapeNet. The feature extractor CTF-Net was replaced by other representative point cloud feature extractors, including PointNet [Qi_2017_CVPR], PointMLP [pointmae], DGCNN [dgcnn], PointMAE [pointmae], Stratified-Transformer [stratifiedtransformer], and PointTransformer-v3 [pointtransformerv3], and some results are reported in Table VII. For transformer-based methods, which are capable of processing all patches in parallel, we adopt the evaluation protocol of MC3D-AD and utilize 10,000 patches to ensure a fair comparison. In contrast, convolution-based baselines and our proposed approach process patches in a sequential manner due to their architectural characteristics.